# Hybrid inflation and gravitational waves from accidentally light scalars

### v1

@@ Acknowledgements
We thank Pierre Auclair, Yann Gouttenoire, Thomas Hambye, Antonio Junior Iovino, Lucas Pinol, Cristophe Ringeval, Simon Schreyer and Alfredo Urbano for useful discussions. MF has received support from the European Union Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreements No 860881-HIDDeN and No 101086085–ASYMMETRY.
@@
Appendix A A model with complex scalars, giving rise to cosmic strings
@@
In Section 2 we presented a minimal model with two real scalar $\mathrm{SU}(2)$ multiplets, a five-plet $\phi$ and a triplet $\chi$. Here we take these fields to be complex, charged under additional $\mathrm{U}(1)$ symmetries.
@@
$V_{\phi}=-\mu_{\phi}^{2}\,\phi^{\dagger}\phi+\lambda\,(\phi^{\dagger}\phi)^{2}+\kappa\left[(\phi^{\dagger}\phi)^{2}-|\phi^{T}\phi|^{2}\right]+\delta\,(\phi^{\dagger}T^{a}\phi)^{2}\,,$ (A1)
with the generators $T^{a}$ defined below Eq. (1). A detailed study of this potential can be found in ref. [3]. [^8] The quartic couplings $\lambda,\kappa,\delta$ here are normalised to be a factor 2 smaller than in ref. [3]. The quartic couplings $\lambda$, $\kappa$ and $\delta$ and the mass parameter $\mu_{\phi}^{2}$ are chosen positive. Notice that the $\kappa$ and $\delta$ terms vanish when $\phi$ is real. Cubic terms are forbidden by $\mathrm{U}(1)_{\phi}$, with no need for ad-hoc discrete symmetries. The minimum of the potential is at $\langle\phi^{\dagger}\phi\rangle=v^{2}/2$, where $v^{2}=\mu_{\phi}^{2}/\lambda$. When $\mathrm{SU}(2)\times\mathrm{U}(1)_{\phi}$ is gauged, all NGBs are absorbed by the gauge bosons. However, there remains a continuous family of vacua at the tree level, parameterized by an angle $\alpha\in[0,\pi/6]$ as in Eq. (2). Thus $\alpha$ is an accidentally flat direction. For generic $\alpha$, all of the gauge symmetry is broken, but at the special point $\alpha=0$, a subgroup $\mathrm{U}(1)^{\prime}\subset\mathrm{SU}(2)$ remains unbroken. At this point there is an additional massless scalar in the spectrum, besides the one corresponding to the flat direction.
@@
Loop corrections lift the accidentally flat direction, analogously to the minimal model of Section 2, with $\alpha=0$ being selected as the true vacuum. The one-loop effective potential along the tree-level flat direction takes the form $V_{\rm eff}\simeq V_{0}^{\prime}-\left(M^{\prime}\right)^{4}\cos\left(a/f\right)$ (up to higher harmonics whose coefficients are numerically small for all values of the couplings of interest). The amplitude of the potential is given by
$\left(M^{\prime}\right)^{4}=\frac{v^{4}}{160\,\pi^{2}}\left[9\,g^{4}-\frac{\kappa^{5}}{\delta^{3}}\left(F\left(-\frac{2\delta}{\kappa}\right)-T_{F,5}\left(-\frac{2\delta}{\kappa}\right)\right)\right]\,,$ (A2)
@@
$\begin{split}V_{\chi}=&\;-\mu^{2}_{\chi}\chi^{\dagger}\chi+\lambda_{\chi}\,(\chi^{\dagger}\chi)^{2}+\lambda^{\prime}_{\chi}\,|\chi^{T}\chi|^{2}\,,\\ V_{\phi\chi}=&\;\varepsilon\,\phi^{\dagger}\phi\chi^{\dagger}\chi+\zeta\,T^{a}_{AB}T^{b}_{BC}\phi^{*}_{A}\phi_{C}\chi^{*a}\chi^{b}+\vartheta\,T^{a}_{AB}(i\varepsilon^{abc})\phi^{*}_{A}\phi_{B}\chi^{*b}\chi^{c}\,.\end{split}$ (A3)
Notice that one can replace the phase rotations $\mathrm{U}(1)_{\phi}\times\mathrm{U}(1)_{\chi}$ by a single gauged $\mathrm{U}(1)$, with charges $q_{\phi}$ and $q_{\chi}$ chosen to forbid additional mixed couplings, beside $\epsilon,\zeta$ and $\vartheta$. In this case, by construction, the scalar potential has an accidental global $\mathrm{U}(1)_{\phi}\times\mathrm{U}(1)_{\chi}$ symmetry. As a consequence, when $\langle\chi\rangle\neq 0$ is generated after inflation, one produces cosmic strings with observable consequences, as discussed in section 5.
For $\mu^{2}_{\chi}<0$ and positive quartic couplings in Eq. (A3), the tree-level potential is easily seen to be minimized at $\chi=0$, with the tree-level vacuum structure for $\phi$ unchanged. In fact, when $\phi$ is set to its tree-level minimum of Eq. (2), the terms $\mu^{2}_{\chi}$ and $\varepsilon$ give rise to a universal, $\alpha$-independent mass term for the three components of $\chi$, while the $\zeta$ term splits their masses in an $\alpha$-dependent fashion (and the $\vartheta$ term does not contribute to the mass matrix). The mass eigenvalues for the $\chi$ fields are the same as in Eq. (4). To obtain a model of hybrid inflation, we choose $\tilde{\mu}_{\chi}^{2}\equiv\mu_{\chi}^{2}-\frac{\varepsilon}{2}v^{2}>0$, and $\zeta v^{2}>2\tilde{\mu}_{\chi}^{2}$, such that the overall $\chi^{3}$ mass-squared is positive for $\alpha=\pi/6$ but negative for $\alpha=0$. At the intermediate value $\alpha=\alpha_{c}$ (given in Eq. (9)) the effective $\chi^{3}$ mass crosses zero.
@@
$\langle\lvert\chi^{3}\rvert\rangle=\sqrt{\frac{\mu^{2}_{\chi}}{2\left(\lambda_{\chi}+\lambda_{\chi}^{\prime}\right)}}\,,\qquad\text{all other fields}=0$ (A4)
@@
Loops of $\chi$ contribute to the accident effective potential so that Eq. (A2) becomes
$\begin{split}\left(M^{\prime}\right)^{4}=\frac{v^{4}}{160\,\pi^{2}}\Bigg{[}&9\,g_{2}^{4}-\frac{\kappa^{5}}{\delta^{3}}\left(F\left(-\frac{2\delta}{\kappa}\right)-T_{F,5}\left(-\frac{2\delta}{\kappa}\right)\right)\\ &+4\,\frac{\tilde{\mu}_{\chi}^{10}}{\zeta^{3}\,v^{10}}\left({\rm Re}\;F\left(\frac{\zeta v^{2}}{\tilde{\mu}_{\chi}^{2}}\right)-T_{F,5}\left(\frac{\zeta v^{2}}{\tilde{\mu}_{\chi}^{2}}\right)\right)\Bigg{]}\,.\end{split}$ (A5)
@@
The calculation of CMB observables in this model is done in the same way as outlined in Section 3, upon replacing $M\,\rightarrow\,M^{\prime}$ and $V_{0}\,\rightarrow\,V_{0}^{\prime}$.
Appendix B A model with disconnected minima, giving rise to unstable domain walls
In this variant of the model, the potential has two physically distinct minima which can be chosen to be near-degenerate. Let $\phi$ be a real five-plet of $\mathrm{SO}(3)$ which is odd under $\mathbb{Z}_{2\phi}:\phi\to-\phi$, and add to the model a pair of triplet waterfall fields $\chi_{R}$ and $\chi_{I}$; we represent them as a single complex field $\chi=(\chi_{R}+i\chi_{I})/\sqrt{2}$, although no $\mathrm{U}(1)$ symmetry is implied. We do however impose a $\mathbb{Z}_{4\chi}$ symmetry acting as $\chi\rightarrow i\chi$. With these fields and symmetries, the most general renormalisable potential reads
$V=$ $-\frac{1}{2}\mu_{\phi}^{2}\phi^{2}+\frac{\lambda_{\phi}}{4}(\phi^{2})^{2}-\mu_{\chi}^{2}\chi^{*}\chi+\lambda_{\chi}\left(\chi^{*}\chi\right)^{2}+\delta\,{\chi^{*}}^{2}\chi^{2}+\dfrac{1}{2}\left(\kappa\,\chi^{2}\chi^{2}+{\rm h.c.}\right)$ (B1)

### v2

@@ Acknowledgements
We thank Pierre Auclair, Yann Gouttenoire, Thomas Hambye, Antonio Junior Iovino, Lucas Pinol, Cristophe Ringeval, Simon Schreyer, and Alfredo Urbano for useful discussions. M. F. has received support from the European Union Horizon 2020 research and innovation program under the Marie Skłodowska-Curie Grant Agreements No 860881-HIDDeN and No. 101086085–ASYMMETRY.
Appendix A Details on the vacuum structure of the model of Sec. 2
In this appendix, we collect some technical details concerning the minimization of the scalar potential in Eq. (1), as well as the computation of the associated one-loop effective potential.
A.1 Tree-level analysis
Eq. (1) is the most general polynomial of degree $\leq 4$ in $\phi$ and $\chi$ which is compatible with $\mathrm{SO}(3)\times\mathbb{Z}_{2}$ invariance (where $\mathbb{Z}_{2}$ acts as $\phi\,\rightarrow\,-\phi$). To see this, note that each term must contain an even number of $\phi$ factors, and that the symmetric products between two $\phi$’s and between two $\chi$’s decompose as $({\mathbf{5}}\otimes{\mathbf{5}})_{\rm sym}={\mathbf{1}}\oplus{\mathbf{5}}\oplus{\mathbf{9}}$ and $({\mathbf{3}}\otimes{\mathbf{3}})_{\rm sym}={\mathbf{1}}\oplus{\mathbf{5}}$, respectively. To form a singlet bilinear, the only possibilities are $(\phi\phi)_{\mathbf{1}}$ and $(\chi\chi)_{\mathbf{1}}$. Cubic $\mathrm{SO}(3)$-invariants can be constructed as $(\phi\phi)_{\mathbf{5}}\phi$ and $(\chi\chi)_{\mathbf{5}}\phi$, but these are forbidden by the $\mathbb{Z}_{2}$ symmetry. Invariant quartic terms are $(\phi\phi)_{\mathbf{1}}(\phi\phi)_{\mathbf{1}}$, $(\phi\phi)_{\mathbf{5}}(\phi\phi)_{\mathbf{5}}$ and $(\phi\phi)_{\mathbf{9}}(\phi\phi)_{\mathbf{9}}$, the latter two of which can be checked to be proportional to the first; $(\chi\chi)_{\mathbf{1}}(\chi\chi)_{\mathbf{1}}$ and $(\chi\chi)_{\mathbf{5}}(\chi\chi)_{\mathbf{5}}$, which are likewise proportional to each other; $(\chi\chi)_{\mathbf{1}}(\phi\phi)_{\mathbf{1}}$ and $(\chi\chi)_{\mathbf{5}}(\phi\phi)_{\mathbf{5}}$. There are, therefore, four independent quartic invariants, which may be represented by the four quartic terms in Eq. (1).
Moreover, the largest continuous symmetry under which Eq. (1) is invariant is indeed $\mathrm{SO}(3)$. We have checked this explicitly, by considering the full $\mathrm{SO}(8)$ symmetry which acts on the $\phi$ and $\chi$ components in the absence of any potential. We verified that only three $\mathrm{SO}(8)$ generators are preserved by $V$, and they form an $\mathrm{SO}(3)$ algebra. This is a nontrivial statement; cf., e.g., the electroweak sector of the Standard Model, where the most general renormalizable potential for the Higgs doublet which is invariant under $\mathrm{SU}(2)\times\mathrm{U}(1)$ gauge transformations turns out to preserve a larger, “custodial” $\mathrm{SO}(4)$ symmetry.
To find the extrema of the tree-level potential of Eq. (1), we will limit ourselves to the case of positive quartic couplings, which is the one of interest for our inflationary model. For definiteness, we use the $5$-plet generators
$\begin{array}[]{c}T^{1}={\scriptsize{\left(\begin{array}[]{ccccc}&&&-i&\\ &&&&-i\\ &&&\sqrt{3}i&\\ i&&-\sqrt{3}i&&\\ &i&&&\end{array}\right)}},\,T^{2}={\scriptsize{\left(\begin{array}[]{ccccc}&-i&&&\\ i&&\sqrt{3}i&&\\ &-\sqrt{3}i&&&\\ &&&&i\\ &&&-i&\end{array}\right)}},\,T^{3}={\scriptsize{\left(\begin{array}[]{ccccc}&&&&-2i\\ &&&-i&\\ &&0&&\\ &i&&&\\ 2i&&&&\end{array}\right)}}\,.\end{array}$ (A1)
[…]
$V=-\frac{1}{2}\mu_{\phi}^{2}\phi^{2}-\frac{1}{2}\mu_{\chi}^{2}\,\chi_{3}^{2}+\frac{\lambda_{\phi}}{4}(\phi^{2})^{2}+\frac{\lambda_{\chi}}{4}\chi_{3}^{4}+\frac{\varepsilon}{4}\phi^{2}\chi_{3}^{2}+\frac{\zeta}{4}\chi_{3}^{2}\,\phi^{\rm T}(T^{3})^{2}\phi\,,$ (A2)
with $(T^{3})^{2}={\rm diag}\,(4,1,0,1,4)$. Hence, the quartic terms are all positive definite, which shows, in particular, that the potential is bounded from below. We have
$\begin{split}\frac{\partial V}{\partial\phi_{A}}=&\;\left[-\mu_{\phi}^{2}+\lambda_{\phi}\,\phi^{2}+\frac{\varepsilon}{2}\chi_{3}^{2}+\frac{\zeta}{2}\chi_{3}^{2}(T^{3})^{2}_{AA}\right]\phi_{A}\,,\\ \frac{\partial V}{\partial\chi_{3}}=&\;\left[-\mu_{\chi}^{2}+\lambda_{\chi}\,\chi_{3}^{2}+\frac{\varepsilon}{2}\phi^{2}+\frac{\zeta}{2}\phi^{\rm T}(T^{3})^{2}\phi\right]\chi_{3}\,.\end{split}$ (A3)
The critical points are, thus, seen to be as follows.
  • $\phi=0$, $\chi=0$.—This is a local minimum if $\mu_{\phi}^{2}<0$ and $\mu_{\chi}^{2}<0$, a local maximum if both are positive (which is the case in our inflationary model), and a saddle point otherwise.
  • $\chi=0$ and $\phi^{2}=\mu_{\phi}^{2}/\lambda_{\phi}\equiv v^{2}$ (assuming $\mu_{\phi}^{2}>0$).— Therefore, this set of critical points forms a four-sphere. The mass matrix does not mix $\phi$ and $\chi$, and the $\phi$ block is easily seen to have eigenvalues $(2\lambda_{\phi}v^{2},0,0,0,0)$. For a generic point on the vacuum manifold, there are three Nambu-Goldstone directions, as expected for a completely broken $\mathrm{SO}(3)$ symmetry, and one additional, accidentally flat direction. However, the ${\bf 5}$-plet generators each have a zero eigenvalue, and, therefore, annihilate a $\phi$ VEV oriented in the direction of the corresponding eigenvector. These special points in the vacuum manifold preserve an O$(2)\subset\mathrm{SO}(3)$ symmetry, and, thus, feature two Nambu-Goldstone fields and two accidentally massless fields. All such points are related to one another by $\mathrm{SO}(3)$ transformations, and, therefore, correspond to a single physical point when $\mathrm{SO}(3)$ is gauged.
    The $\chi$ block of the mass matrix is
    $\left.\frac{\partial^{2}V}{\partial\chi_{a}\partial\chi_{b}}\right|_{\chi=0,\,\phi^{2}=v^{2}}=\left(\frac{\varepsilon}{2}v^{2}-\mu_{\chi}^{2}\right)\delta^{ab}+\frac{\zeta}{2}\phi_{A}T^{a}_{AC}T^{b}_{CB}\phi_{B}\,.$ (A4)
[…]
  • $\phi=0$ and $\chi^{2}=\mu_{\chi}^{2}/\lambda_{\chi}\equiv v_{\chi}^{2}$ (assuming $\mu_{\chi}^{2}>0$).—The mass matrix in the $\chi$ sector has eigenvalues $(2\lambda_{\chi}v_{\chi}^{2},0,0)$, with two Nambu-Goldstone directions as expected for $\mathrm{SO}(3)\,\rightarrow\,\mathrm{SO}(2)$ breaking by a triplet VEV. The mass matrix in the $\phi$ sector is diagonal in a basis where $\chi_{1,2}=0$ and $\chi_{3}=v_{\chi}$:
    $\left.\frac{\partial^{2}V}{\partial\phi_{A}\partial\phi_{B}}\right|_{\chi=\begin{pmatrix}0\\ 0\\ v_{\chi}\end{pmatrix},\,\phi=0}=\bigg{[}\frac{\varepsilon}{2}v_{\chi}^{2}-\mu_{\phi}^{2}+\frac{\zeta}{2}v_{\chi}^{2}(T^{3})^{2}_{AA}\bigg{]}\delta_{AB}\,.$ (A5)
    Therefore, this critical point is a minimum provided that $\frac{\varepsilon}{2}v_{\chi}^{2}>\mu_{\phi}^{2}$. It is the true vacuum of our inflationary model, by a choice of parameter values.
[…]
A.2 One-loop corrections
At the one-loop level, the tree-level scalar potential is corrected by the well-known Coleman-Weinberg effective potential
$V_{\rm eff}=\frac{1}{64\pi^{2}}\,\operatorname{tr}\,\left({\cal M}^{4}\log\frac{{\cal M}^{2}}{\Lambda^{2}}\right)\,.$ (A6)
Here, ${\cal M}$ is the scalar-field-dependent mass matrix for gauge fields and scalars (with a factor 3 for the former, accounting for three polarizations for a massive vector field) and $\Lambda$ is the renormalization scale.
We are interested, notably, in the loop-corrected effective potential for a field configuration with VEVs $\phi^{2}=v^{2}$ and $\chi=0$, along the direction which remained accidentally flat at the tree level. This direction was parameterized by an angle $\alpha\in[0,\,\pi/6]$ in the main text; see Eq. (2). Evaluating $V_{\rm eff}$ at some fixed value of $\alpha$, $\bar{\alpha}=\pi/6$, say, we find that there is a divergent tadpole term for the radial mode $\rho$, as well as a divergent contribution to the vacuum energy density. These divergences need to be subtracted by including counterterms for all parameters of the model and imposing suitable renormalization conditions. A particularly convenient set of conditions is to (i) renormalize the $\rho$ tadpole to be zero, such that $v^{2}\equiv m_{\phi}^{2}/\lambda_{\phi}$ remains the VEV after including one-loop corrections (this is an implicit condition on the counterterms for $m^{2}_{\phi}$ and $\lambda_{\phi}$); and (ii) set the renormalized vacuum energy density to some finite constant. There are other divergences in the one-loop effective action, but these two conditions suffice to render the effective potential along the $\alpha$ direction finite. It is given by [3]
$\begin{array}[]{l}V_{\rm eff}(\alpha)=\dfrac{1}{64\pi^{2}}\times\\ \times\Biggl{[}\operatorname{tr}\left(M^{4}(\alpha)\log\dfrac{M^{2}(\alpha)}{\Lambda^{2}}-M^{4}(\bar{\alpha})\log\dfrac{M^{2}(\bar{\alpha})}{\Lambda^{2}}\right)\\ +3\,\operatorname{tr}\left(M_{W}^{4}(\alpha)\log\dfrac{M_{W}^{2}(\alpha)}{\Lambda^{2}}-M_{W}^{4}(\bar{\alpha})\log\dfrac{M_{W}^{2}(\bar{\alpha})}{\Lambda^{2}}\right)\Biggr{]}\,,\end{array}$ (A7)
where $M$ and $M_{W}$ are the scalar and gauge boson mass matrices, respectively, and the $\Lambda$ dependence is, in fact, spurious and cancels out, $\operatorname{tr}M^{4}$ and $\operatorname{tr}M_{W}^{4}$ being $\alpha$-independent. Note that the choice of the subtraction point $\alpha=\bar{\alpha}=\pi/6$ is arbitrary: We could have chosen any other point provided that the tree-level mass matrix is positive definite, so that the effective potential is well-defined. The gauge symmetry leads to the identifications $\alpha\,\rightarrow\,-\alpha$ and $\alpha\,\rightarrow\,\alpha+\pi/3$; therefore, the effective potential is even and periodic. As such, $V_{\rm eff}(\alpha)$ can be expanded into a Fourier cosine series, $V_{\rm eff}(\alpha)=\sum_{n}c_{n}\cos(6n\alpha)$. The leading Fourier mode of $V_{\rm eff}$ is then found to be given by the expressions of Eqs. (6) and (7), and it provides a slow-roll potential for the inflaton field.
@@ A.2 One-loop corrections
Appendix B A model with complex scalars, giving rise to cosmic strings
@@
In Sec. 2, we presented a minimal model with two real scalar $\mathrm{SU}(2)$ multiplets, a five-plet $\phi$ and a triplet $\chi$. Here, we take these fields to be complex, charged under additional $\mathrm{U}(1)$ symmetries.
@@
$V_{\phi}=-\mu_{\phi}^{2}\,\phi^{\dagger}\phi+\lambda\,(\phi^{\dagger}\phi)^{2}+\kappa\left[(\phi^{\dagger}\phi)^{2}-|\phi^{T}\phi|^{2}\right]+\delta\,(\phi^{\dagger}T^{a}\phi)^{2}\,,$ (B1)
with the generators $T^{a}$ defined below Eq. (1). A detailed study of this potential can be found in Ref. [3]. [^8] The quartic couplings $\lambda,\kappa$, and $\delta$ here are normalized to be a factor 2 smaller than in Ref. [3]. The quartic couplings $\lambda$, $\kappa$, and $\delta$ and the mass parameter $\mu_{\phi}^{2}$ are chosen positive. Notice that the $\kappa$ and $\delta$ terms vanish when $\phi$ is real. Cubic terms are forbidden by $\mathrm{U}(1)_{\phi}$, with no need for ad-hoc discrete symmetries. The minimum of the potential is at $\langle\phi^{\dagger}\phi\rangle=v^{2}/2$, where $v^{2}=\mu_{\phi}^{2}/\lambda$. When $\mathrm{SU}(2)\times\mathrm{U}(1)_{\phi}$ is gauged, all NGBs are absorbed by the gauge bosons. However, there remains a continuous family of vacua at the tree level, parameterized by an angle $\alpha\in[0,\pi/6]$ as in Eq. (2). Thus, $\alpha$ is an accidentally flat direction. For generic $\alpha$, all of the gauge symmetry is broken, but, at the special point $\alpha=0$, a subgroup $\mathrm{U}(1)^{\prime}\subset\mathrm{SU}(2)$ remains unbroken. At this point, there is an additional massless scalar in the spectrum, besides the one corresponding to the flat direction.
@@
Loop corrections lift the accidentally flat direction, analogously to the minimal model in Sec. 2, with $\alpha=0$ being selected as the true vacuum. The one-loop effective potential along the tree-level flat direction takes the form $V_{\rm eff}\simeq V_{0}^{\prime}-\left(M^{\prime}\right)^{4}\cos\left(a/f\right)$ (up to higher harmonics whose coefficients are numerically small for all values of the couplings of interest). The amplitude of the potential is given by
$\left(M^{\prime}\right)^{4}=\frac{v^{4}}{160\,\pi^{2}}\Bigg{[}9\,g^{4}-\frac{\kappa^{5}}{\delta^{3}}\Bigg{(}F\left(-\frac{2\delta}{\kappa}\right)-T_{F,5}\left(-\frac{2\delta}{\kappa}\right)\Bigg{)}\Bigg{]}\,,$ (B2)
@@
$V_{\chi}=$ $\;-\mu^{2}_{\chi}\chi^{\dagger}\chi+\lambda_{\chi}\,(\chi^{\dagger}\chi)^{2}+\lambda^{\prime}_{\chi}\,|\chi^{T}\chi|^{2}\,,$ (B3)
$V_{\phi\chi}=$ $\;\varepsilon\,\phi^{\dagger}\phi\chi^{\dagger}\chi+\zeta\,T^{a}_{AB}T^{b}_{BC}\phi^{*}_{A}\phi_{C}\chi^{*a}\chi^{b}+\vartheta\,T^{a}_{AB}(i\varepsilon^{abc})\phi^{*}_{A}\phi_{B}\chi^{*b}\chi^{c}\,.$
Notice that one can replace the phase rotations $\mathrm{U}(1)_{\phi}\times\mathrm{U}(1)_{\chi}$ by a single gauged $\mathrm{U}(1)$, with charges $q_{\phi}$ and $q_{\chi}$ chosen to forbid additional mixed couplings, beside $\epsilon,\zeta$, and $\vartheta$. In this case, by construction, the scalar potential has an accidental global $\mathrm{U}(1)_{\phi}\times\mathrm{U}(1)_{\chi}$ symmetry. As a consequence, when $\langle\chi\rangle\neq 0$ is generated after inflation, one produces cosmic strings with observable consequences, as discussed in Sec. 5.
For $\mu^{2}_{\chi}<0$ and positive quartic couplings in Eq. (B3), the tree-level potential is easily seen to be minimized at $\chi=0$, with the tree-level vacuum structure for $\phi$ unchanged. In fact, when $\phi$ is set to its tree-level minimum of Eq. (2), the terms $\mu^{2}_{\chi}$ and $\varepsilon$ give rise to a universal, $\alpha$-independent mass term for the three components of $\chi$, while the $\zeta$ term splits their masses in an $\alpha$-dependent fashion (and the $\vartheta$ term does not contribute to the mass matrix). The mass eigenvalues for the $\chi$ fields are the same as in Eq. (4). To obtain a model of hybrid inflation, we choose $\tilde{\mu}_{\chi}^{2}\equiv\mu_{\chi}^{2}-\frac{\varepsilon}{2}v^{2}>0$, and $\zeta v^{2}>2\tilde{\mu}_{\chi}^{2}$, such that the overall $\chi^{3}$ mass-squared is positive for $\alpha=\pi/6$ but negative for $\alpha=0$. At the intermediate value $\alpha=\alpha_{c}$ (given in Eq. (9)), the effective $\chi^{3}$ mass crosses zero.
@@
$\langle\lvert\chi^{3}\rvert\rangle=\sqrt{\frac{\mu^{2}_{\chi}}{2\left(\lambda_{\chi}+\lambda_{\chi}^{\prime}\right)}}\,,\qquad\text{all other fields}=0$ (B4)
@@
Loops of $\chi$ contribute to the accident effective potential so that Eq. (B2) becomes
$\begin{split}\left(M^{\prime}\right)^{4}=\frac{v^{4}}{160\,\pi^{2}}\Bigg{[}&9\,g_{2}^{4}-\frac{\kappa^{5}}{\delta^{3}}\left(F\left(-\frac{2\delta}{\kappa}\right)-T_{F,5}\left(-\frac{2\delta}{\kappa}\right)\right)\\ &+4\,\frac{\tilde{\mu}_{\chi}^{10}}{\zeta^{3}\,v^{10}}\left({\rm Re}\;F\left(\frac{\zeta v^{2}}{\tilde{\mu}_{\chi}^{2}}\right)-T_{F,5}\left(\frac{\zeta v^{2}}{\tilde{\mu}_{\chi}^{2}}\right)\right)\Bigg{]}\,.\end{split}$ (B5)
@@
The calculation of CMB observables in this model is done in the same way as outlined in Sec. 3, upon replacing $M\,\rightarrow\,M^{\prime}$ and $V_{0}\,\rightarrow\,V_{0}^{\prime}$.
Appendix C A model with disconnected minima, giving rise to unstable domain walls
In this variant of the model, the potential has two physically distinct minima which can be chosen to be near-degenerate. Let $\phi$ be a real five-plet of $\mathrm{SO}(3)$ which is odd under $\mathbb{Z}_{2\phi}:\phi\to-\phi$, and add to the model a pair of triplet waterfall fields $\chi_{R}$ and $\chi_{I}$; we represent them as a single complex field $\chi=(\chi_{R}+i\chi_{I})/\sqrt{2}$, although no $\mathrm{U}(1)$ symmetry is implied. We do, however, impose a $\mathbb{Z}_{4\chi}$ symmetry acting as $\chi\rightarrow i\chi$. With these fields and symmetries, the most general renormalizable potential reads
$V=$ $-\frac{1}{2}\mu_{\phi}^{2}\phi^{2}+\frac{\lambda_{\phi}}{4}(\phi^{2})^{2}-\mu_{\chi}^{2}\chi^{*}\chi+\lambda_{\chi}\left(\chi^{*}\chi\right)^{2}+\delta\,{\chi^{*}}^{2}\chi^{2}+\dfrac{1}{2}\left(\kappa\,\chi^{2}\chi^{2}+{\rm h.c.}\right)$ (C1)